the joint distribution function of X and Y is given by
F(x,y) = (1-e-x)(1-e-y) for x>0,y>0
= 0 otherwise
Discuss about the marginal densities of X and Y
Examine whether X and Y are independent (iii) Evaluate P(1<x<3,1<y<2)
1
Expert's answer
2020-04-13T18:44:25-0400

1) Marginal pdf of X:

"f_X(x)=\int_{-\infty}^{\infty}f_{X, Y}(x,y)dy.\\
f_{X, Y}(x,y)\text{ --- joint pdf}.\\
f_{X, Y}(x,y)=\frac{\partial^2}{\partial x\partial y}F(x,y).\\
F(x,y)\text{ --- joint CDF}.\\
\frac{\partial}{\partial x}F(x,y)=(1-e^{-y})e^{-x}.\\
\frac{\partial^2}{\partial x\partial y}F(x,y)=e^{-(x+y)}.\\
f_{X, Y}(x,y)=e^{-(x+y)} (x>0, y>0); 0 \text{ (elsewhere)}.\\
f_X(x)=\int_{0}^{\infty}e^{-(x+y)}dy=e^{-x}.\\
\text{So marginal pdf of X }f_X(x)=e^{-x} (x>0); 0 \text{ (elsewhere)}.\\
\text{Marginal pdf of Y (all the computations are symmetrical)}\\
f_Y(y)=e^{-y} (y>0); 0 \text{ (elsewhere)}.\\
2) X\text{ and } Y\text{ are independent because } f_{X,Y}(x,y)=f_X(x)f_Y(y).\\
P\{1<x<3, 1<y<2\}=\int_{1}^{3}\int_{1}^{2}f_{X,Y}(x,y)dydx=\int_{1}^{3}\int_{1}^{2}f_{X}(x)f_{Y}(y)dydx=\int_{1}^{3}\int_{1}^{2}e^{-x}e^{-y}dydx=(e^{-2}-e^{-1})(e^{-3}-e^{-1})\approx\\
\approx 0.074."